a pair of dice is thrown 3 times. if getting 10 as a sum is success. find probability distribution of success.

Grade:12

1 Answers

Vijay Mukati
8 years ago
Dear Student,

Let X denotes the number of times we get sucess, So the possible values of X will be {0, 1, 2, 3}

Now Probability of sucess = 3/36 = 1/12 for each trial (5+5 = 6+4 = 4+6).
so probability of faliure = 11/12

P(X=0) = (11/12)*(11/12)*(11/12)
P(X=1) = 3C1*(1/12)*(11/12)*(11/12)
P(X=2) = 3C2*(1/12)*(1/12)*(11/12)
P(X=3) = (1/12)*(1/12)*(1/12)

Thanks.